OpenStudy (anonymous):

Alegbra Help Anyone? D': -4(x + 1) - 3 = -3(x - 4)

OpenStudy (johnweldon1993):

Hi again :) Step 1: Distribute the -4 into the parenthesis...and do the same with that -3

OpenStudy (anonymous):

i dont understand distributing

OpenStudy (anonymous):

you have to multiply -4 to x and also to 1

OpenStudy (johnweldon1993):

Distributing...you take that -4...and you multiply it by every term inside the parenthesis... -4(x + 1) would become (-4 * x) + (-4 * 1) simplifying a bit more. -4x - 4 make sense?

OpenStudy (anonymous):

yeah i think, sorry im like really really bad at math /.\

OpenStudy (anonymous):

do i do it to the -3(x-4) too?

OpenStudy (johnweldon1993):

No one is bad at math...you may have just never learned it in a good way...just like that last problem I helped you on....your teachers didnt explain it in a way you understood :)

OpenStudy (johnweldon1993):

and yes...do it to that as well

OpenStudy (anonymous):

then what about the other -3?

OpenStudy (johnweldon1993):

That'll come later...okay so what do you get when you distribute that -3(x - 4) ?

OpenStudy (anonymous):

-12x?

OpenStudy (johnweldon1993):

Not quite...remember you cannot combine a number with a variable to another number. -3(x - 4) (-3 * x) - (-3 * 4) -3x + 12 right?

OpenStudy (anonymous):

mkay (: now what /.\

OpenStudy (johnweldon1993):

Now it's easy... From everything we did...we have left... -4x - 4 - 3 = -3x + 12 ^ theres that -3 Okay...now it is just combining everything...but like I said...we can only combine numbers with an 'x'...with another number with an 'x' -4x and -3x can be combined.. every other term is just a number....which can all be combined with each other... So simplify that whole equation..what do you get?

OpenStudy (anonymous):

so i take -4x and add it too -3x? then add everything else?

OpenStudy (johnweldon1993):

Yeah you can do that

OpenStudy (anonymous):

-7 + 5 ?

OpenStudy (johnweldon1993):

Not exactly...so it simplifies like this -4x - 4 - 3 = -3x + 12 you want to bring all the 'x' to 1 side...and all the other numbers to the other side So lets add 3x to both sides of this equation... -4x - 4 - 3 = -3x + 12 +3x +3x ------------------------ -x - 4 - 3 = 12 Make sense? we an only have 1 x ... Now..we want to isolate that 'x'....lets move ALL the other numbers to the other side of the eqaution... -x - 4 - 3 = 12 Well lets add 4 and add 3 to both sides... -x - 4 - 3 = 12 + 4 + 3 +4 + 3 ----------------------- -x = 19 Right? now the ultimate goal is to solve for 'x' what does 'x' equal...notice it is 'x' not '-x' so we have 1 more step...and that is to switch the signs of the 2 terms we have -x = 19 will become x = -19 See how x became positive and 19 became negative... so our final answer here...will be x = -19

OpenStudy (anonymous):

thank you.!